# Global Non-Axisymmetric Hall Instabilities in a Rotating Plasma

### v1

@@ Abstract
Non-axisymmetric, flow-driven instabilities in the incompressible Hall-MHD model are studied in a differentially rotating cylindrical plasma. It is found that in the Hall-MHD regime, both whistler waves and ion-cyclotron waves can extract energy from the flow shear, resulting in two distinct branches of global instability. The non-axisymmetric whistler modes grow significantly faster than non-axisymmetric, ideal MHD modes. A discussion of the whistler instability mechanism is presented in the large-ion-skin-depth, ‘electron-MHD’ limit. It is observed that the effect of the Hall term on the non-axisymmetric modes can be appreciable when $d_{i}$ is on the order of a few % of the width of the cylindrical annulus. Distinct global modes emerge in the Hall-MHD regime at significantly stronger magnetic fields than those required for unstable global MHD modes.
@@ 1 Introduction
Weakly magnetized, differentially rotating plasmas are unstable to flow-driven hydromagnetic instabilities (E. Velikhov, 1959; S. Chandrasekhar, 1960). The relevance of this mechanism to accretion disks was explored in the seminal work of S. A. Balbus & J. F. Hawley (1991), wherein the phenomenon was given the name magneto-rotational instability (MRI). The study of such linear instabilities is motivated by the question of whether their onset leads to the generation of turbulence, and if the associated turbulent viscosity contributes significantly rate of angular momentum transport in accretion disks (N. I. Shakura & R. A. Sunyaev, 1973). While the MRI descriptions of E. Velikhov (1959), S. Chandrasekhar (1960) and S. A. Balbus & J. F. Hawley (1991) use ideal magnetohydrodynamics (MHD) to model the ionized gases in an accretion disk, it has been noted that plasma behavior beyond ideal MHD is likely important in weakly ionized protoplanetary disks (PPDs) (O. M. Blaes & S. A. Balbus, 1994). Other work (M. Wardle, 1999; S. A. Balbus & C. Terquem, 2001; M. W. Kunz & S. A. Balbus, 2004; F. Ebrahimi et al., 2011) has specifically considered the influence of the Hall term, $\boldsymbol{J}\times\boldsymbol{B}/en_{e}$, on the linear dispersion relation for the axisymmetric MRI using local, WKB-like approximations of the governing differential equations. Linear studies of MRI modes in non-ideal regime have shown that inclusion of the Hall term modifies the axisymmetric dispersion relation in two ways. First, S. A. Balbus & C. Terquem (2001) note that Hall effect allows instability in systems that have a radially increasing rotation rate. Such rotation profiles are predicted to be stable in MHD limit. The second observation is that the local dispersion relation becomes asymmetric with respect to the orientation of the magnetic field relative to the axis of rotation/vorticity of the flow. Local linear theory was developed by M. W. Kunz (2008) for a non-rotating flow instability in Hall-MHD using a shearing box approximation that was dubbed the Hall-shear instability (HSI). The flow shear driven Hall-MHD instability in non-rotating plasmas has also been referred to as a magnetoshear instability (MSI) (C. Bejarano et al., 2011). 3D Hall-MHD calculations show that the MSI can out-compete the growth of Kelvin-Helmholtz instabilities in a shear layer with perpendicular magnetic field if the vorticity is sufficiently large (D. O. Gómez et al., 2014). It has also been shown that there exists a second branch of instability in the local dispersion Hall-MHD dispersion relation that results from coupling of the ion-cyclotron wave to the epicyclic motion of a differentially rotating disc (J. B. Simon et al., 2015). In these local linear calculations (M. Wardle, 1999; S. A. Balbus & C. Terquem, 2001; M. W. Kunz, 2008; C. Bejarano et al., 2011; J. B. Simon et al., 2015), it was shown that growth rate depends on the relative orientation of the vorticity and the magnetic field, and instability is possible for either positive or negative shear (increasing or decreasing rotation rate).
@@
Recent work solving the radial eigenvalue problem has elucidated the importance of global non-axisymmetric modes in the MHD model due to the cylindrical curvature of the geometry and flow curvature (F. Ebrahimi & M. Pharr, 2022; F. Ebrahimi & A. Haywood, 2025; A. Haywood & F. Ebrahimi, 2025a), as well as in laboratory settings (Y. Wang et al., 2025). Alfvénic resonances (R. Matsumoto & T. Tajima, 1995; G. I. Ogilvie & J. E. Pringle, 1996; F. Ebrahimi & M. Pharr, 2022) are shown to be instrumental in the global nature of the MHD non-axisymmetric flow-driven and curvature modes (the global non-axisymmetric MRI branch). The question arises whether dispersive Hall-MHD modes contribute to the onset of global flow-driven instabilities in differentially rotating systems. Here, we uncover that in fact whistler and ion-cyclotron waves can have a dominant effect on the onset of non-axisymmetric modes in the weakly-ionzied regime (Hall-MHD model).
@@
This work presents a comprehensive global linear study of non-axisymmetric, flow driven instabilities in the Hall MHD model. In the Hall-MHD model, we show that there are two distinct branches of instability in the presence of the Hall effect. One branch is faster growing the MHD MRI modes at modest values of the ion skin depth, $d_{i}$. We identify this branch as a whistler wave that is able to extract energy from the steady-state flow shear. The second branch is akin to the incompressible MHD modes with growth rates and mode structures that are modified by the hall term. We also explore the instabilities of the rotating system in the large $d_{i}$, electron-MHD (EMHD) limit. In this regime, the ions provide an effectively stationary, neutralizing background through which the electron fluid moves. This model retains the coupling between the mean rotation profile and the whistler waves. In this way, the electron MHD model shows that there is a distinct branch of linear instability that is not simply explained as a destabilizing effect on the non-axisymmetric MRI modes. Both local and global analysis and calculations of the electron MHD system are presented.
@@
Using a combination of local and global eigenvalue analysis as well as numerical calculations, we assess the degree to which the Hall-MHD waves extract energy from the differential rotation - leading to new non-axisymmetric modes that exist at larger magnetic field strengths than MHD modes (F. Ebrahimi & M. Pharr, 2022). We find that for $k=0$ modes in the EMHD model, the local instability criterion predicts stability despite the presence of global unstable whistler modes.
@@
The organization of this paper is as follows. Section 2 presents analysis and numerical calculations of the linear, non-axisymmetric instabilities of the rotating disk in the EMHD limit. It is shown that non-axisymmetric whistler waves are driven unstable by the sheared rotation when the wave frequency is similar in magnitude to the rotation rate. The local, point-wise dispersion relation theory is compared to the results of global analysis. Necessary conditions for the existence of unstable modes are derived based on global properties of the EMHD system. Numerical results for the global eigenvalue problem are presented at several values of disk aspect ratio (height to width), in an attempt to study the role of the boundary locations on global instabilities. It is found that lower aspect ratio (thinner disk) generally leads to a instability over a broader range of magnetic field strengths. Section 3 investigates the unstable modes in the full Hall-MHD system for both magnetic field configurations. The strength of the hall term is varied by adjusting the ion skin depth, $d_{i}=c/\omega_{pi}$, where $\omega_{pi}\equiv\sqrt{Z^{2}e^{2}n_{i}/\epsilon_{0}m_{i}}$ is the ion plasma frequency.
1.1 Basic Equations
The incompressible Hall-MHD equations for a fully-ionized plasma linearized about a steady, differentially rotating azimuthal flow are given, in cylindrical coordinates, by
@@ 1.1 Basic Equations
$\Omega$ is the azimuthal rotation frequency, $\kappa^{2}\equiv 4\Omega^{2}-2\Omega r\Omega^{\prime}$ is the epicyclic frequency, $\boldsymbol{B}$ is the background magnetic field, and $d_{i}=c/\omega_{pi}$ is the ion skin depth. The lowercase variables $\boldsymbol{v},\boldsymbol{b}$ represent the perturbed flow and magnetic field respectively, and $\tilde{P}=p+\boldsymbol{B}\cdot\boldsymbol{b}/\mu_{0}$ is the perturbed total pressure. We note that the identification $1/en_{e}\rightarrow d_{i}/\sqrt{\mu_{0}\rho}$ is only valid to lowest order in $m_{e}/m_{i}$, and assumes $\rho\approx n_{e}m_{i}$. If one were to consider instead, a weakly-ionized plasma where $\rho$ and $\boldsymbol{v}$ correspond to the total bulk density and flow including neutral gas, $d_{i}$ should to be replaced by $\ell_{H}\equiv\sqrt{(\rho/\rho_{i})}d_{i}$ (B. P. Pandey & M. Wardle, 2008).
The axial field is presumed uniform throughout the domain, and the azimuthal field satisfies the current-free condition; $\partial_{r}(rB_{\phi})=0$. Following S. Chandrasekhar (1960) and E. Frieman & M. Rotenberg (1960), it is convenient to introduce the Lagrangian displacement $\boldsymbol{\xi}$ such that $\boldsymbol{v}=\partial_{t}\boldsymbol{\xi}+\boldsymbol{V}\cdot\nabla\boldsymbol{\xi}-\boldsymbol{\xi}\cdot\nabla\boldsymbol{V}$. In terms of $\boldsymbol{\xi}$, the system can be written as an eigenvalue problem for a Hermitian differential system whose eigenvalues are $\omega^{2}$ when $m=0,k\neq 0$ (S. Chandrasekhar, 1961). In that case, the problem of determining stability of the system is reduced to finding parameters for which solutions of the system of ODEs satisfying the boundary conditions have negative eigenvalues. In the present case, $m\neq 0$, we elect to write the equations in terms of $\boldsymbol{\xi}$ for the sake of comparison with prior results.
@@
Here, $\bar{\omega}\equiv\omega-m\Omega(r)$, $F\equiv\boldsymbol{k}\cdot\boldsymbol{B}$, $\omega_{A}\equiv F/\sqrt{\mu_{0}\rho}$, and $V_{H}\equiv\omega_{A}d_{i}$. $\omega_{A}$ is the local frequency of a shear Alfvén wave, and $V_{H}$ is the phase velocity of a parallel propagating whistler wave. Compared to the linear, incompressible, ideal MHD equations, equations (4)-(7) include two additional first-order differential equations in the radial coordinate. In particular, we note that the radial components of the perturbed displacement and magnetic field are no longer simply proportional to each other. This reflects the fact that in non-dissipative Hall-MHD, the magnetic field is ‘frozen in’ to the electron fluid motion instead of the bulk ion motion (E. Hameiri et al., 2005). This is in contrast to the ideal MHD case, where eq. (6) simplifies to $\psi=F\chi$.
@@
The dispersion relation for a magnetized, homogeneous plasma in the incompressible Hall-MHD limit possesses two sets of roots: $(\omega^{2}-\omega_{A}^{2})^{2}=k^{2}d_{i}^{2}\omega^{2}\omega_{A}^{2}$. The two solutions of this dispersion relation are plotted as a function of $kd_{i}$ in figure 1. The pair of solutions with the larger phase velocity describe whistler waves when $kd_{i}$ is large, and are analogous to the MHD fast magnetosonic wave in the compressible case at low-$kd_{i}$. The slower solution for $\omega^{2}$ is the Hall-modified shear Alfvén wave in the low-$kd_{i}$ regime, which transitions into ion cyclotron waves propagating parallel to $\boldsymbol{B}$ in the large $kd_{i}$ limit [^1] In the incompressible case, the slow wave and shear Alfvén wave are degenerate. The shear Alfvén mode is incompressible at low $kd_{i}$, but the slow wave branch is incompressible at large $kd_{i}$. (E. Hameiri et al., 2005). In the incompressible limit, the fast wave and the shear Alfvén wave are degnerate at $kd_{i}\ll 1$ (see figure 1). However, as we will show in section 3, small but finite $d_{i}$ lifts the degeneracy between the two wave branches, leading to qualitatively different behavior in the differentially rotating flow.
@@
Figure 1: Plot of the positive solutions of $(\omega^{2}-\omega_{A}^{2})=\pm kd_{i}\omega\omega_{A}$ as a function of $kd_{i}$. $\Omega_{ci}\equiv|V_{A}|/d_{i}$ is the ion cyclotron frequency, and $\omega_{A}\equiv k_{\parallel}B/\sqrt{\mu_{0}\rho}$ is the Alfvén frequency. Here, $k_{\parallel}=k\cos(\pi/6)$

### v2

@@ Abstract
Non-axisymmetric, flow-driven instabilities in the incompressible Hall-MHD model are studied in a differentially rotating cylindrical plasma. It is found that in the Hall-MHD regime, both whistler waves and ion-cyclotron waves can extract energy from the flow shear, resulting in two distinct branches of global instability. The non-axisymmetric whistler modes grow significantly faster than non-axisymmetric, ideal MHD modes. A discussion of the global whistler instability mechanism is presented in the large-ion-skin-depth, ‘electron-MHD’ limit. When the magnetic field is azimuthal, a subset of the whistler modes having zero axial wave number are uncovered to be destabilized by the ‘co-rotation amplifier’ mechanism. It is observed that the effect of the Hall term on the non-axisymmetric modes can be appreciable when $d_{i}$ is on the order of a few % of the width of the cylindrical annulus. Distinct global modes emerge in the strong Hall-MHD regime at significantly stronger magnetic fields than those required for unstable global MHD modes, as the Hall effect weakens the stabilizing ‘field-line bending’ by decoupling ion motion from the magnetic field. These global non-axisymmetric modes may play an important role in weakly ionized accretion disks.
@@ I Introduction
Weakly magnetized, differentially rotating plasmas are unstable to flow-driven hydromagnetic instabilities (Velikhov, 1959; Chandrasekhar, 1960). The relevance of this mechanism to accretion disks was explored in the seminal work of Balbus and Hawley (1991), wherein the phenomenon was given the name magnetorotational instability (MRI). The study of these linear instabilities is motivated by the question of whether their onset leads to the generation of turbulence, and if the associated turbulent viscosity contributes significantly to the rate of angular momentum transport in accretion disks that is required by $\alpha$-disk models (Shakura and Sunyaev, 1973). While the MRI descriptions of Velikhov (1959), Chandrasekhar (1960) and Balbus and Hawley (1991) use ideal magnetohydrodynamics (MHD) to model the ionized gases in an accretion disk, it has been noted that plasma behavior beyond ideal MHD is likely important in weakly ionized protoplanetary disks (PPDs) (Blaes and Balbus, 1994). Other work (Wardle, 1999; Balbus and Terquem, 2001; Kunz and Balbus, 2004; Ebrahimi et al., 2011) has specifically considered the influence of the Hall term, $\boldsymbol{J}\times\boldsymbol{B}/en_{e}$, on the linear dispersion relation for the axisymmetric MRI using local, WKB-like approximations of the governing differential equations. Linear studies of MRI modes in the non-ideal regime have shown that inclusion of the Hall term modifies the axisymmetric dispersion relation in two ways. First, Balbus and Terquem (2001) note that Hall effect allows instability in systems that have a radially increasing rotation rate. Increasing rotation profiles are predicted to be stable in ideal MHD. The second observation is that the local dispersion relation becomes asymmetric with respect to the orientation of the magnetic field relative to the axis of rotation/vorticity of the flow. Local linear theory was developed by Kunz (2008) for a non-rotating flow instability in Hall-MHD using a shearing box approximation that was dubbed the Hall-shear instability (HSI). The flow shear driven Hall-MHD instability in non-rotating plasmas has also been referred to as a magnetoshear instability (MSI) (Bejarano et al., 2011). 3D Hall-MHD calculations show that the MSI can out-compete the growth of Kelvin-Helmholtz instabilities in a shear layer with perpendicular magnetic field if the vorticity is sufficiently large (Gómez et al., 2014). It has also been shown that there exists a second branch of instability in the local Hall-MHD dispersion relation that results from coupling of the ion-cyclotron wave to the epicyclic motion of a differentially rotating disc (Simon et al., 2015). In these local linear calculations (Wardle, 1999; Balbus and Terquem, 2001; Kunz, 2008; Bejarano et al., 2011; Simon et al., 2015), it was shown that growth rate depends on the relative orientation of the vorticity and the magnetic field, and instability is possible for either positive or negative shear (increasing or decreasing rotation rate).
@@
Recent work solving the radial eigenvalue problem has elucidated the importance of global non-axisymmetric modes in the MHD model due to the cylindrical curvature of the geometry and flow curvature (Ebrahimi and Pharr, 2022; Ebrahimi and Haywood, 2025; Haywood and Ebrahimi, 2025), as well as in laboratory settings (Wang et al., 2025). Alfvénic resonances (Matsumoto and Tajima, 1995; Ogilvie and Pringle, 1996; Ebrahimi and Pharr, 2022) are shown to be instrumental in the global nature of the MHD non-axisymmetric flow-driven and curvature modes (the global non-axisymmetric MRI branch). The question arises whether dispersive Hall-MHD modes contribute to the onset of global flow-driven instabilities in differentially rotating systems. Here, we uncover that in fact whistler and ion-cyclotron waves can have a dominant effect on the onset of non-axisymmetric modes in the Hall-MHD model.
@@
This work presents a linear study of the spectrum of non-axisymmetric, flow driven instabilities in the Hall-MHD model in a rotating plasma Couette flow of the type one might expect near the mid-plane of a protoplanetary accretion disk. In the Hall-MHD model, we show that there are two distinct branches of instability in the presence of the Hall effect. One branch is faster growing than the MHD MRI modes at modest values of the ion skin depth, $d_{i}$. We identify this branch as a whistler wave that is able to extract energy from the steady-state flow shear. The second branch is akin to the incompressible MHD modes with growth rates and mode structures that are modified by the Hall term. We also explore the instabilities of the rotating system in the large $d_{i}$, electron-MHD (EMHD) limit. In this regime, ions provide an effectively stationary, neutralizing background through which the electron fluid moves. This model retains the coupling between the mean rotation profile and the whistler waves. In this way, the EMHD model isolates the whistler branch of linear instability that is not simply explained as a destabilizing effect on the non-axisymmetric MHD MRI modes. Both local and global analysis and numerical calculations of the electron MHD system are presented.
@@
Using a combination of local and global eigenvalue analysis as well as numerical calculations, we assess the degree to which the Hall-MHD waves extract energy from the differential rotation - leading to new non-axisymmetric modes that exist at larger magnetic field strengths than MHD modes (Ebrahimi and Pharr, 2022). We find that for purely parallel propagating modes in the EMHD model, the local WKB theory predicts stability despite the presence of global unstable whistler modes.
@@
The organization of this paper is as follows. Section II introduces the Hall-MHD model and investigates incompressible, non-axisymmetric instabilities in differentially-rotating cylindrical disk geometry. We present numerical results using two different methods. The strength of the hall term is varied by adjusting the ion skin depth, $d_{i}=c/\omega_{pi}$, where $\omega_{pi}\equiv\sqrt{Z^{2}e^{2}n_{i}/\epsilon_{0}m_{i}}$ is the ion plasma frequency. Section III presents analysis and numerical calculations of the EMHD limit. It is shown that non-axisymmetric whistler waves are driven unstable by the sheared rotation when the wave frequency is similar in magnitude to the rotation rate. The local, point-wise dispersion relation theory is compared to the results of global analysis. Necessary conditions for the existence of unstable modes are derived based on global properties of the EMHD system. Numerical results for the global eigenvalue problem are presented at several values of disk aspect ratio (height to width), in an attempt to study the role of the boundary locations on global whistler instabilities. It is found that lower aspect ratio (thinner disk) generally leads to instability over a broader range of magnetic field strengths.
II The Hall-MHD model
We consider an incompressible Hall-MHD model of plasma dynamics. The full system of nonlinear PDEs couples the bulk fluid motion of a quasi-neutral two-component plasma to ‘pre-Maxwell’ classical electromagnetism. The equations describing the fluid motions include an equation of continuity for the plasma mass density, $\rho$, moving with velocity field $\boldsymbol{v}$.
$\partial_{t}\rho+\nabla\cdot\left(\rho\boldsymbol{v}\right)=0.$ (1)
Assuming electron inertia is negligible, the equation governing the total momentum evolution is
$\rho\partial_{t}\boldsymbol{v}+\rho\boldsymbol{v}\cdot\nabla\boldsymbol{v}=\frac{\boldsymbol{B}\cdot\nabla\boldsymbol{B}}{\mu_{0}}-\nabla P-\rho\nabla\Phi_{g}.$ (2)
Here, $P$ is the sum of the isotropic gas pressure, $p$, and the magnetic pressure, $B^{2}/2\mu_{0}$. $\Phi_{g}$ is the Newtonian gravitational potential. We assume the source of the gravitational potential is a central body of mass $M$, and neglect self-gravitation of the disk material, so that $\Phi_{g}=-GM/r$. The time-dependence of the magnetic field $\boldsymbol{B}$ is determined by Faraday’s law,
$\partial_{t}\boldsymbol{B}=-\nabla\times\boldsymbol{E},$ (3)
and the electric field is given by the Hall-MHD Ohm’s law,
$\boldsymbol{E}=-\boldsymbol{v}\times\boldsymbol{B}+\frac{\boldsymbol{J}\times\boldsymbol{B}}{en_{e}}.$ (4)
Here, $n_{e}$ is the electron density, which can be approximately related to the mass density, $\rho$ by $n_{e}\sim Z\rho/m_{i}$ for a two-component plasma with ions with charge $Ze$, and $m_{e}\ll m_{i}$. In the MHD approximation, we neglect displacement current, so $\mu_{0}\boldsymbol{J}=\nabla\times\boldsymbol{B}$. The system of equations is completed by incompressibility, $\nabla\cdot\boldsymbol{v}=0$, and the magnetic divergence constraint $\nabla\cdot\boldsymbol{B}=0$.
To address the question of linear stability, the equations are linearized around a steady-state, equilibrium solution in cylindrical $(r,\phi,z)$ coordinates (see figure 1). The steady-state is defined by an axisymmetric, rotating, uniform-density, current-free azimuthal flow ($\boldsymbol{v}=r\Omega\hat{\boldsymbol{\phi}}$) that is supported by either pressure gradients or a gravitational potential. The cylindrical domain is intended to represent an annular slice of the accreting material very near the mid-plane of the disk. We assume that the vertical domain is small enough that we may neglect stratification of the plasma and the vertical components of $\nabla\Phi_{g}$. We also assume azimuthal symmetry of the steady-state. Thus, variation of the steady-state fields in the direction away from the disk mid-plane and in the azimuthal direction is neglected; $\partial_{\phi}Q=\partial_{z}Q=0$ for any $Q\in\{\Omega,\boldsymbol{B},\Phi_{g},P,\rho\}$. The remaining non-trivial relations for the steady-state are
$\partial_{r}B_{z}=\partial_{r}(rB_{\phi})=0,$ (5)
and
$\rho r\Omega^{2}=\partial_{r}P+\rho\frac{GM}{r^{2}}.$ (6)
If $\partial_{r}P=0$, $\Omega=\sqrt{GM/r^{3}}$ is the Keplerian rotation rate – the orbital frequency of circular orbits in the gravitational potential created by the central body. Arbitrary $\Omega(r)$ profiles can be accommodated in the steady-state by including a nonzero radial pressure gradient. In the incompressible case considered in this work, the linearized equations are unaffected by the steady-state radial pressure gradient.
Figure 1: Schematic representation of the cylindrical rotating plasma equilibrium.
To study linear stability of the steady flow, we linearize the equations by perturbing each quantity by a small amount. The magnetic field, for example, becomes $\boldsymbol{B}+\boldsymbol{b}$. Here, $\boldsymbol{B}$ is a solution to the steady state described above, and we assume $|\boldsymbol{b}|/|\boldsymbol{B}|\ll 1$. Applying this procedure to each variable in equations (1)-(4), and neglecting any terms that depend on higher powers of perturbed quantities, one arrives at the following incompressible linearized Hall-MHD equations for a fully-ionized plasma linearized about a steady, differentially rotating azimuthal flow are given, in cylindrical coordinates, by
@@ II The Hall-MHD model
$\kappa^{2}\equiv 4\Omega^{2}+2\Omega r\Omega^{\prime}$ is the epicyclic frequency, $\boldsymbol{B}$ is the background magnetic field, and $d_{i}=c/\omega_{pi}$ is the ion skin depth. The lowercase variables $\boldsymbol{v},\boldsymbol{b}$ represent the perturbed flow and magnetic field respectively, and $\tilde{P}=p+\boldsymbol{B}\cdot\boldsymbol{b}/\mu_{0}$ is the perturbed total pressure. We note that the identification $1/en_{e}\rightarrow d_{i}/\sqrt{\mu_{0}\rho}$ is only valid to lowest order in $m_{e}/m_{i}$, and assumes $\rho\approx n_{e}m_{i}$. If one were to consider a weakly-ionized plasma where $\rho$ and $\boldsymbol{v}$ correspond to the total bulk density and flow including neutral gas, then $d_{i}$ ought to be replaced by $\ell_{H}\equiv\sqrt{(\rho/\rho_{i})}d_{i}$ (Pandey and Wardle, 2008).
The axial component of the magnetic field, $B_{z}$, is presumed uniform throughout the domain, and the azimuthal component satisfies the current-free condition; $\partial_{r}(rB_{\phi})=0$. No assumption is made regarding the specific form of $\Omega(r)$ while deriving the following equations. However, since it is anticipated than Hall effects are most dominant in protoplanetary accretion disks (Wardle and Salmeron, 2012), we will consider the case of Keplerian rotation where $\Omega\propto r^{-3/2}$ when computing numerical values for the growth rates.
Following the work of Chandrasekhar (1960) and Frieman and Rotenberg (1960) in the ideal MHD case, it is convenient to introduce the Lagrangian displacement $\boldsymbol{\xi}$ such that $\boldsymbol{v}=\partial_{t}\boldsymbol{\xi}+\boldsymbol{V}\cdot\nabla\boldsymbol{\xi}-\boldsymbol{\xi}\cdot\nabla\boldsymbol{V}$. The ideal MHD system can be rewritten as a single second-order ODE in the radial component of the displacement $\xi_{r}$ (Khalzov et al., 2006; Ebrahimi and Pharr, 2022; Ebrahimi and Haywood, 2025). If the perturbations are axisymmetric, $\partial_{\phi}\boldsymbol{\xi}=\partial_{\phi}{\boldsymbol{b}}=0$, the resulting ODE is an eigenvalue problem for a Hermitian differential system whose eigenvalues are $\omega^{2}$ (Chandrasekhar, 1961). In that case, the problem of determining stability of the system is reduced to finding parameters for which a solution of the system of ODEs satisfying given boundary conditions has a negative eigenvalue. In the present case we do not assume axisymmetry, so the problem cannot be reduced to finding the eigenvalues of a Hermitian differential operator. Furthermore, the inclusion of the Hall term complicates the system so that reduction to a single second order equation in $\xi_{r}$ is not possible. However, we elect to write the equations in terms of $\boldsymbol{\xi}$ for the sake of comparison.
@@
Here, $\bar{\omega}\equiv\omega-m\Omega(r)$, $F\equiv\boldsymbol{k}\cdot\boldsymbol{B}$, $\omega_{A}\equiv F/\sqrt{\mu_{0}\rho}$, and $V_{H}\equiv\omega_{A}d_{i}$. $\omega_{A}$ is the local frequency of a shear Alfvén wave, and $V_{H}$ is the phase velocity of a parallel propagating whistler wave. Compared to the linear, incompressible, ideal MHD equations, equations (10)-(13) include two additional first-order differential equations in the radial coordinate that complicate the relationship between the magnetic field and fluid motion. In particular, we note that the radial components of the perturbed displacement and magnetic field are no longer simply proportional to each other. The additional terms in equations (12) and (13) reflect the fact that in non-dissipative Hall-MHD, the magnetic field is ‘frozen in’ to the electron fluid motion instead of the bulk ion motion (Hameiri et al., 2005). This is in contrast to the ideal MHD case, where equation (12) simplifies to $\psi=-F\chi$.
@@
To orient our discussion of the Hall effect on the spectrum of instabilities present in the system, it is useful to first consider the dispersion relation for a magnetized, homogeneous plasma in the incompressible Hall-MHD limit. It possesses two sets of roots: $(\omega^{2}-\omega_{A}^{2})^{2}=k^{2}d_{i}^{2}\omega^{2}\omega_{A}^{2}$. The two solutions of this dispersion relation are plotted as a function of $kd_{i}$ in figure 2. The pair of solutions with the larger phase velocity describe whistler waves when $kd_{i}$ is large, and are analogous to the MHD fast magnetosonic wave in the compressible case at low-$kd_{i}$. The slower solution for $\omega^{2}$ is the Hall-modified shear Alfvén wave in the low-$kd_{i}$ regime, which transitions into ion cyclotron waves propagating parallel to $\boldsymbol{B}$ in the large $kd_{i}$ limit [^1] In the incompressible case, the slow wave and shear Alfvén wave are degenerate. The shear Alfvén mode is incompressible at low $kd_{i}$, but the slow wave branch is incompressible at large $kd_{i}$. (Hameiri et al., 2005). In the incompressible limit, the fast wave and the shear Alfvén wave are degnerate at $kd_{i}\ll 1$ (see figure 2). However, as we will show in Section II.1, small but non-zero $d_{i}$ lifts the degeneracy between the two wave branches, leading to qualitatively different behavior in the differentially rotating flow.
@@
Figure 2: Plot of the positive solutions of $(\omega^{2}-\omega_{A}^{2})=\pm kd_{i}\omega\omega_{A}$ as a function of $kd_{i}$. $\Omega_{ci}\equiv|V_{A}|/d_{i}$ is the ion cyclotron frequency, $V_{A}=B/\sqrt{\mu_{0}\rho}$ is the Alfvén velocity and $\omega_{A}\equiv k_{\parallel}B/\sqrt{\mu_{0}\rho}$ is the associated frequency of shear-Alfvén waves. Here, we choose nearly parallel propagation: $k_{\parallel}=k\cos(\pi/6)$